OpenStudy (anonymous):

The numbers on two consecutively numbered gym lockers have a sum of 155. What are the locker numbers?

OpenStudy (campbell_st):

well let x be the 1st number then the next consecutive locker is x + 1 so you know the sum of the 2 lockers then x + x + 1 = 155 now solve for x

OpenStudy (anonymous):

x^+1=155

OpenStudy (anonymous):

right?

OpenStudy (anonymous):

x+x = 2x Then remember. that will find you the first number. your second number will be that answer plus one.

OpenStudy (anonymous):

wait so huh sorry I had to d something. so is it just 77 or more

OpenStudy (anonymous):

your one answer will be 77. that is your x. now you have to find x+1. (77+1) Your 2 answers will be 155 when you add them together.

OpenStudy (anonymous):

232 right?

OpenStudy (anonymous):

am I right or wrong?

OpenStudy (anonymous):

No, it's asking you to find 2 numbers. They are in order, so for instance, 2, 3. Together they have to add up to 155. We decided the first number was going to be called x. The second number will be called x+1. You found x to be 77. That is your first number, so that is the number of the first locker. The second locker will then be one more than the first, so 78. You can check this by putting your numbers into your equation. x+x+1=155 77+77+1=155

OpenStudy (anonymous):

so 77 and 78

OpenStudy (anonymous):

yes. do you see why?

OpenStudy (anonymous):

yes they both add up to 155 thaks a lot

OpenStudy (anonymous):

no problem